OpenStudy (anonymous):

Point L is located at (4, -3) and M is located at (-8, 5). Find the coordinates of the point that lies halfway between L and M. Select one: a. (0.5, -1.5) b. (6, -4) c. (3.5, -6.5) d. (-2, 1)

OpenStudy (anonymous):

@ganeshie8 @satellite73

OpenStudy (anonymous):

Use this mid-point formula \[ (x_m,y_m) = \left ( \frac { x_1+x_2}2, \frac { y_1+y_2}2 \right) \]

OpenStudy (anonymous):

\[ (x_m,y_m) = \left ( \frac {4+(-8)}2, \frac { -3+5}2\right) =(-2,1) \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw